OpenStudy (anonymous):

Help!!!! Simplify 4|cos θ| if θ = sin−1 (x/4) for some real number x. isn't the answer √16-x^2???

OpenStudy (paxpolaris):

 the answer is\[\sqrt{16-x^2}\]... as long as \(-4\le x\le 4\)

OpenStudy (paxpolaris):

you are right

OpenStudy (anonymous):

it doesn't say the range..

OpenStudy (anonymous):

you are sooooooo awesome!!!